distance between point A & point B = 310 KM . two person leave these points at the same time , one with speed of 106km/h and the other with 87km/h . How long does it take to meet each other ?
Answer by gonzo(654)

AB = 310 km.
AC = x
BC = 310-x
-----
T = time
-----
person leaving from point A is traveling at 87km/h.
person leaving from point B is traveling at 106km/h.
-----
they both meet at point C.
-----
person leaving from point A travels x kilometers in T hours.
person leaving from point B travels (310-x) kilometers in T hours.
-----
rate * time = distance
-----
person leaving from A equation:
87*T = x
-----
person leaving from B equation:
106*T = 315-x
-----
solve both equations for T
-----
T = x/87 (first equation)
T = (315-x)/106 (second equation)
-----
since they both equal T then they both equal each other.
x/87 = (315-x)/106
-----
multiply both sides by (87*106)
equation becomes
106*x = (315-x)*87
remove parentheses.
106*x = 315*87 - 87*x
add 87*x to both sides.
106*x + 87*x = 315*87
perform operations shown.
193*x = 27405
divide both sides by 193.
x = 141.9948187
solve for T in first equation.
x/87 = T
141.9948187/87 = T
T = 1.632124352
-----
315-x = 173.0051813
-----
calculate T from second equation.
T = (315-x)/106
T = 173.0051813/106
T = 1.632124352
-----
point C is 141.9948187 miles from A
point C is 173.0051813 miles from B
AC + BC = 315.
-----
rate of travel equals distance over time.
141.9948187 / 1.632124352 = 87 km/h
173.0051813 / 1.632124352 = 106 km/h
-----
answer is T = 1.632124352 hours
They will meet at point C in 1.632124352 hours